OpenStudy (anonymous):

hich of the following is the solution to the equation 363c = square root of 6(6c + 8

OpenStudy (anonymous):

\[363c=\sqrt{6(6c+8)}\rightarrow363^{2}c ^{2}=36c+48\]

OpenStudy (anonymous):

my choices are -3/2 -4/3 4/3 and 3/2

OpenStudy (anonymous):

\[363^{2}c ^{2}-36c-48=0\] \[c=(36\pm \sqrt{36^{2}-4*363^{2}*48})/(2*363^{2})\] just calculate this

OpenStudy (anonymous):

sry, under the root should be +, not -

OpenStudy (anonymous):

i dont have a calculator handy can u do it please